40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 5 09:20:44 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_Chamber_optical_layout.pdf
    Reply  Thu Aug 5 12:39:56 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_in-air_setup.pdf
       Reply  Thu Aug 5 14:29:42 2021, Stephen, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber IMG_9685.JPGIMG_9687.JPGrouting_markup_of_IMG_9688.pngIMG_9682.JPG
          Reply  Tue Aug 24 15:01:25 2021, Stephen, Laser, 2um Photodiodes, 2um PD in new cryo chamber 
             Reply  Thu Aug 26 14:31:46 2021, Aidan, Stephen, Laser, 2um Photodiodes, Incident power calibration of 2um PD in new cryo chamber IMG_4420.jpgREF_PD_POW_CAL_cryo.pdf
                Reply  Fri Aug 27 17:57:30 2021, Stephen, DailyProgress, 2um Photodiodes, Start of PD pumpdown  
                   Reply  Sat Aug 28 07:01:21 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  20210828_-_IMG_4445.jpg20210828_-_IMG_4443.jpg
                      Reply  Sun Aug 29 11:17:43 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  Screen_Shot_2021-08-29_at_11.16.55_AM.pngScreen_Shot_2021-08-29_at_11.22.03_AM.png
                         Reply  Mon Aug 30 08:04:52 2021, Aidan, DailyProgress, 2um Photodiodes, Monday morning observation of JPL A1 PD cool down  Screen_Shot_2021-08-30_at_8.04.15_AM.png
Message ID: 2651     Entry time: Thu Aug 26 14:31:46 2021     In reply to: 2650     Reply to this: 2654
Author: Aidan, Stephen 
Type: Laser 
Category: 2um Photodiodes 
Subject: Incident power calibration of 2um PD in new cryo chamber 

We're getting close to running this (2um beam is focussed onto the PD and we have piezo mirror steering and beam size tuning). However, all DAC channels are currently non-responsive. I'm going to rebooting the front-end.

Incident power calibration was performed with a S148C power meter head placed directly in front of the PD. For varying current levels, I recorded the pick-off PD voltage and the power meter reading.

Current (mA) REF PD (V) Power Meter (mW)
101.04 1.01 0.803
101.03 1.012 0.805
100.41 1.01 0.805
93.11 0.935 0.735
85.03 0.846 0.664
79.97 0.792 0.618
75.01 0.74 0.573
69.97 0.684 0.524
65 0.629 0.478
55.01 0.52 0.384
44.96 0.405 0.284
35 0.291 0.1834
25 0.176 0.0844
20 0.119 0.0348
0.03 0.079 3.00E-04

POW [mW] = 0.862[mW/V]*REF_PD [V] -0.067[mW]

Quote:

[Aidan, Stephen]

Worked toward aligning and characterizing beam on PD. Will complete next session.

Log:

Some difficulty aligning to the 2um beam, which is sensed by a thermal card. Aidan intends to upgrade with a fiber coupled visible laser, which could then be swapped interchangably for alignment.

The 1" mirror at the top of the periscope doesn't make sense, given larger apertures in shields and viewport. We looked for a nearby 2" replacement but did not have luck. Ended up swapping back in the gold-coated 2" mirror, even though it is thin enough to be a pain to mount.

Instrumented connector pins to DB9 pins using the following translation (ref Aidan's drawing for connector / PD pinout, ref drawing from QIL/2639)

            DB9   - 1 6 2 7 3 8 

    Connector - 2 3 4 5 6 7

 

 

Attachment 1: IMG_4420.jpg  871 kB  Uploaded Thu Aug 26 17:05:54 2021  | Hide | Hide all
IMG_4420.jpg
Attachment 2: REF_PD_POW_CAL_cryo.pdf  8 kB  Uploaded Fri Aug 27 08:48:11 2021  | Hide | Hide all
REF_PD_POW_CAL_cryo.pdf
ELOG V3.1.3-